OpenStudy (anonymous):

18y=z, solve the equation for y

OpenStudy (anonymous):

\[y = z/18\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

18y=z, solve the equation for y [divide both sides by 18] y = z/18

OpenStudy (anonymous):

y=z/18

OpenStudy (anonymous):

Thank you